### v1

@@ Abstract
High-dimensional sparse data emerge in many critical application domains such as cybersecurity, healthcare, anomaly detection, and trend analysis. To quickly extract meaningful insights from massive volumes of these multi-dimensional data, scientists employ unsupervised analysis tools based on tensor decomposition (TD) methods. However, real-world sparse tensors exhibit highly irregular shapes, data distributions, and sparsity, which pose significant challenges for making efficient use of modern parallel architectures. Moreover, as the number of cores per processor continues to increase, memory contention and synchronization become substantial performance bottlenecks for irregular workloads. This study breaks the prevailing assumption that compressing sparse tensors into coarse-grained structures (i.e., tensor slices or blocks) or along a particular dimension/mode (i.e., mode-specific) is more efficient than keeping them in a fine-grained, mode-agnostic form. Our novel sparse tensor representation, Adaptive Linearized Tensor Order (ALTO), encodes tensors in a compact format that can be easily streamed from memory and is amenable to both caching and parallel execution. In contrast to existing compressed tensor formats, ALTO constructs one tensor copy that is agnostic to both the mode orientation and the irregular distribution of nonzero elements. To demonstrate the efficacy of ALTO, we accelerate popular TD methods that compute the Canonical Polyadic Decomposition (CPD) model across a range of real-world sparse tensors. We propose a set of parallel TD algorithms that exploit the inherent data reuse of tensor computations to substantially reduce synchronization overhead, decrease memory footprint, and improve parallel performance. Additionally, we characterize the major execution bottlenecks of TD methods on multiple generations of the latest Intel Xeon Scalable processors, including Sapphire Rapids CPUs, and introduce dynamic adaptation heuristics to automatically select the best algorithm based on the sparse tensor characteristics. Across a diverse set of real-world data sets, ALTO outperforms the state-of-the-art approaches, achieving more than an order-of-magnitude speedup over the best mode-agnostic formats. Compared to the best mode-specific formats, which require multiple tensor copies, ALTO achieves more than geometric mean speedup at a fraction () of their storage.
@@
Sparse tensors, tensor decomposition, MTTKRP, Poisson tensor decomposition, multi-core CPU, Sapphire Rapids1 Introduction
@@
Tensors, the higher-order generalization of matrices, can naturally represent complex interrelations in multi-dimensional sparse data, which emerge in important application domains such as healthcare [1, 2], cybersecurity [3, 4], data mining [5, 6], and machine learning [7, 8]. For example, one mode (or dimension) of a tensor may identify users while another mode details their demographic information, and their (potentially incomplete) ratings for a set of products [9]. To effectively analyze such high-dimensional data, tensor decomposition (TD) is used to reveal their principal components, where each component represents a latent property. One of the most popular TD models is the Canonical Polyadic Decomposition (CPD), which approximates a tensor as a sum of a finite number of rank-one tensors such that each rank-one tensor corresponds to a tensor component, or a latent property [10, 11]. An important class of real-world, high-dimensional data sets is non-negative sparse tensors with count data [12, 13], which encodes critical information, such as the number of packets exchanged across a network, or the number of criminal activities in a city. For these tensors, the CP Alternating Poisson Regression (CP-APR) algorithm is a powerful tool for detecting anomalies and group relations. In contrast to other CPD algorithms that assume a Gaussian distribution for randomly distributed data (e.g., CP Alternating Least Squares, or CP-ALS), CP-APR [14] assumes a Poisson distribution which better describes the target count data.
@@
The curse of dimensionality causes data to become more dispersed as the number of modes increases. Hence, high-dimensional data sets typically suffer from highly irregular shapes and data distributions as well as unstructured and extreme sparsity, which make them challenging to represent efficiently. For instance, Figure 1 illustrates the spatial distribution of nonzero elements in a set of sparse tensors. It shows that the number of nonzero elements in a subspace, or a multi-dimensional block, can vary greatly (note the use of logarithmic scale). Furthermore, as the sparsity of tensors increases (e.g., deli, nell-1, amazon, and reddit tensors), the likelihood of finding dense structures in the multi-dimensional space significantly decreases, leading to extremely small numbers of nonzero elements per block. Therefore, efficient execution of TD algorithms on modern parallel architectures is challenging because of their low arithmetic intensity [15], random memory access, workload imbalance [16, 17], and data dependencies [15, 17]. Moreover, TD algorithms typically require computations along every mode orientation, and realizing acceptable performance across all modes is difficult without using multiple mode-specific tensor copies.
@@
Prior work on the CP-ALS algorithm improved the parallel performance of the matricized tensor times Khatri-Rao product (MTTKRP) kernel [15, 16, 18, 20, 21], which is the main performance bottleneck and can make up as much as 90% of the execution time of the overall algorithm [20]. The few performance studies conducted on the CP-APR algorithm focused primarily on performance portability [22] and streaming analysis [23], rather than parallel performance optimization. To the best of our knowledge, this paper presents the first in-depth analysis of the key performance bottlenecks of CP-APR, and significantly improves the parallel performance of both CP-ALS and CP-APR over prior state-of-the-art approaches by using a linearized mode-agnostic sparse tensor representation.
@@
Additionally, the previous approaches relied on extending legacy sparse linear algebra formats and algorithms to tensor (multilinear) algebra problems [11, 24, 20, 25, 26, 27, 18, 28, 29, 30]. These techniques can be classified based on their compression of the nonzero elements into raw or compressed formats [31, 32]. Raw formats use list-based tensor representations to keep the nonzero elements along with their multi-dimensional coordinates [33]. Hence, they are mode-agnostic and typically require a single tensor copy to execute tensor operations along different modes. As a result, the list-based coordinate (COO) format remain the de facto data structure for storing sparse tensors [33] in many tensor libraries (e.g., Tensor Toolbox [11], Tensorflow [34], and Tensorlab [35]). However, due to their unprocessed nature, list-based formats suffer from significant parallel and synchronization overheads [26].
@@
Compressed tensor formats [20, 36, 18, 28] use tree- or block-based structures to organize the nonzero elements, which may decrease the memory footprint of sparse tensors. However, since these approaches rely on finding clusters of nonzero elements in non-overlapping regions of the multi-dimensional space to achieve compression, their efficacy depends on the spatial data distributions, which can be highly irregular and extremely sparse in higher-order data sets as demonstrated in Figure 1. Therefore, instead of reducing memory storage, compressed formats can introduce substantial memory overhead and degrade the parallel performance and scalability of TD algorithms [31].
@@
The most popular compressed format for TD algorithms is compressed sparse fiber (CSF) [20], which extends the classical compressed sparse row (CSR) format to higher-order tensors using tree-like structures. However, CSF-based formats are mode-specific, where the arrangement of nonzero elements depends not only on the mode being considered as the root of the index tree but also on the other modes at each subsequent tree level; therefore, they are typically efficient for only that specific mode order. This leads to a trade-off between performance and memory, as storing multiple copies of the tensor (i.e., copies for a mode- tensor), each arranged for a specific mode, yields the best performance at the cost of extra memory [36]. For large-scale tensors, keeping multiple copies may be infeasible, especially for hardware accelerators with limited memory capacity (e.g., GPUs). In contrast, keeping only a single copy of the tensor, arranged for one arbitrarily chosen mode order, yields the smallest memory footprint at the cost of sub-optimal performance [25]. This performance degradation can be significant depending on how the nonzero elements are distributed in the -dimensional space [31], and traversing the nonzero elements for computing along different modes requires entirely different algorithms [25]. While block-based compressed formats [18, 28] can be mode-agnostic, their storage requirements and parallel performance still depend on the spatial data distributions as well as the parameters of the blocking/tilling schemes [28, 31], which are difficult to determine dynamically.
@@
To overcome these limitations, we present the Adaptive Linearized Tensor Order (ALTO) format. ALTO is a mode-agnostic sparse representation that maps a set of -dimensional coordinates onto a single linearized index such that neighboring nonzero elements in the multi-dimensional space are close to each other in memory, which results in a more cache friendly and memory-scalable tensor storage; that is, ALTO utilizes the inherent data locality of sparse tensors and its storage scales with mode lengths, rather than the number of modes. Additionally, ALTO enables a unified implementation of tensor algorithms that requires a single tensor copy to compute along all modes.
@@
In contrast to prior compressed [20, 18] and linearized [37] TD approaches, we propose a set of parallel sparse tensor algorithms that leverage the ALTO format to address the performance bottlenecks that have traditionally limited the scalability of irregular tensor computations. Specifically, our proposed parallel TD algorithms generate perfectly balanced tensor partitions in terms of the number of nonzero elements; however, these partitions may divide the multi-dimensional space of a tensor into overlapping regions. Thus, we present data-aware adaptation heuristics that greatly reduce synchronization overhead and improve parallel performance by locating the overlapping space between tensor partitions and dynamically selecting the best tensor traversal and conflict resolution method according to the inherent data locality of sparse tensors. Moreover, these heuristics choose between recomputing or reusing intermediate results, depending on the properties of the target tensor computations, to maximize the performance while reducing the overall memory footprint. As a result, our ALTO-based TD algorithms deliver substantial performance improvements over prior state-of-the-art approaches, while allowing the processing of large-scale tensors and the computing of high decomposition ranks. In summary, we make the following contributions:
@@
We present ALTO, a novel sparse tensor format for high-performance and scalable execution of tensor operations. Unlike prior compressed sparse tensor formats, ALTO uses a single (mode-agnostic) tensor representation that improves data locality, eliminates workload imbalance, and greatly reduces memory usage and synchronization overhead, regardless of the data distribution in the multi-dimensional space (3).
@@
We propose efficient ALTO-based parallel algorithms for the CP-ALS and CP-APR methods as well as input-aware adaptation heuristics for parallel conflict resolution and memory management (4).
@@
We conduct an in-depth performance analysis of the main TD kernels and compare our parallel algorithms against prior state-of-the-art across two generations of the latest Intel Xeon Scalable processors. The results demonstrate that ALTO-based tensor algorithms achieve and geometric mean speedup over the best mode-agnostic and mode-specific formats, respectively, while realizing geometric mean compression over the best mode-specific formats (5).
@@ 2.1 Tensor Notations
Tensors are -dimensional arrays, where each element has a corresponding -tuple index . Each index coordinate locates a tensor element along the dimension or mode, with and . Low-dimensional tensors include vectors, where , and matrices, where . In general, a dense -dimensional (or a mode-) tensor has indexed elements. A tensor is said to be sparse if the majority of its elements are zero. The following notations are used in this paper:
@@ 2.2.1 CP-ALS
Algorithm 1 illustrates the widely used Alternating Least Squares (ALS) method for iteratively computing the factor matrices of the CPD model. In each CP-ALS iteration, every factor matrix is updated to yield the best approximation of when all other factor matrices are fixed. Line 13 shows the matricized tensor times Khatri-Rao product (MTTKRP) operation [20], which involves tensor matricization and Khatri-Rao product, and it is typically the most expensive tensor kernel of the CP-ALS algorithm. For a mode- tensor , the mode-1 MTTKRP operation can be expressed as . Typically, MTTKRP operations along all modes are performed 10–100 times in a single tensor decomposition calculation. Since these MTTKRP operations are similar across all modes, for brevity, we only discuss mode- MTTKRP in this paper.
@@ 2.2.2 CP-APR
While CP-ALS can decompose a sparse count tensor, CP-APR better describes the random variations in the data by representing it using a Poisson distribution [14], which considers a discrete number of events and assumes zero probability for observing fewer than zero events. Yet, CP-APR is more expensive to compute compared to CP-ALS [14]. There are three common methods for computing CP-APR: (i) Multiplicative update (MU), (ii) Projected damped Newton for row-based sub-problems (PDN-R), and (iii) Projected quasi-Newton for row-based sub-problems (PQN-R).
@@
The CP-APR algorithm using MU is shown in Algorithm 2. Using two nested loops, CP-APR computes the decomposition model of a tensor by successively updating each factor matrix while holding the other factor matrices fixed. Lines 8 and 10 show the (Khatri-Rao product) and (model update) kernels, respectively, which make up majority of the CP-APR execution time. Note that is calculated once per outer loop for each mode, and is calculated once per inner loop (the inner loop is executed at most times per mode per outer loop), and the cost of calculating can more easily be amortized as the number of inner iterations goes up.
@@ 2.3.1 Coordinate (COO)
Decomposing a sparse tensor stored in the COO format typically involves iterating over each nonzero element and updating the corresponding factor matrix row. For example, during mode- computation, a nonzero element with coordinates (,,), updates row of the mode- factor matrix after reading rows and from mode- and mode- factor matrices, respectively. Since multiple threads can simultaneously update the same row of a factor matrix, these updates must be done atomically, which can be expensive on parallel processors with a large number of threads.

### v2

@@ Abstract
High-dimensional sparse data emerge in many critical application domains such as healthcare and cybersecurity. To extract meaningful insights from massive volumes of these multi-dimensional data, scientists employ unsupervised analysis tools based on tensor decomposition (TD) methods. However, real-world sparse tensors exhibit highly irregular shapes and data distributions, which pose significant challenges for making efficient use of modern parallel processors. This study breaks the prevailing assumption that compressing sparse tensors into coarse-grained structures (i.e., tensor slices or blocks) or along a particular dimension/mode (i.e., mode-specific) is more efficient than keeping them in a fine-grained, mode-agnostic form. Our novel sparse tensor representation, Adaptive Linearized Tensor Order (ALTO), encodes tensors in a compact format that can be easily streamed from memory and is amenable to both caching and parallel execution. In contrast to existing compressed tensor formats, ALTO constructs one tensor copy that is agnostic to both the mode orientation and the irregular distribution of nonzero elements. To demonstrate the efficacy of ALTO, we accelerate popular TD methods that compute the Canonical Polyadic Decomposition (CPD) model across different types of sparse tensors. We propose a set of parallel TD algorithms that exploit the inherent data reuse of tensor computations to substantially reduce synchronization overhead, decrease memory footprint, and improve parallel performance. Additionally, we characterize the major execution bottlenecks of TD methods on multiple generations of the latest Intel Xeon Scalable processors, including Sapphire Rapids CPUs, and introduce dynamic adaptation heuristics to automatically select the best algorithm based on the sparse tensor characteristics. Across a diverse set of real-world data sets, ALTO outperforms the state-of-the-art approaches, achieving more than an order-of-magnitude speedup over the best mode-agnostic formats. Compared to the best mode-specific formats, which require multiple tensor copies, ALTO achieves geometric mean speedup at a fraction () of their storage costs. Moreover, ALTO obtains geometric mean speedup over the state-of-the-art memoization approach, which reduces computations by using extra memory, while requiring of its memory consumption.
@@
Sparse tensors, tensor decomposition, ALTO, MTTKRP, CP-APR, Poisson tensor decomposition, Alternating Poisson Regression, Multi-core CPU, Sapphire Rapids1 Introduction
@@
Tensors, the higher-order generalization of matrices, can naturally represent complex interrelations in multi-dimensional sparse data, which emerge in important application domains such as healthcare [1, 2], cybersecurity [3, 4], data mining [5, 6], and machine learning [7, 8]. For example, one mode (or dimension) of a tensor may identify users while another mode details their demographic information, and their (potentially incomplete) ratings for a set of products [9]. To effectively analyze such high-dimensional data, tensor decomposition (TD) is used to reveal their principal components, where each component represents a latent property. One of the most popular TD models is the Canonical Polyadic Decomposition (CPD), which approximates a tensor as a sum of a finite number of rank-one tensors such that each rank-one tensor corresponds to a tensor component, or a latent property [10, 11]. An important class of real-world, high-dimensional data sets is sparse tensors with non-negative count data [12, 13], which encodes critical information, such as the number of packets exchanged across a network, or the number of criminal activities in a city. For these tensors, the CP Alternating Poisson Regression (CP-APR) algorithm is a powerful tool for detecting anomalies and group relations. In contrast to other CPD algorithms that assume a Gaussian distribution for randomly distributed data (e.g., CP Alternating Least Squares, or CP-ALS), CP-APR [14] assumes a Poisson distribution which better describes the target count data.
@@
Due to the curse of dimensionality, data become more dispersed as the number of modes increases. Hence, high-dimensional data sets typically suffer from highly irregular shapes and data distributions as well as unstructured and extreme sparsity, which make them challenging to represent efficiently. For instance, Figure 1 illustrates the spatial distribution of nonzero elements in a set of sparse tensors. It shows that the number of nonzero elements in a subspace, or a multi-dimensional block, can vary greatly (note the use of logarithmic scale). Furthermore, as the sparsity of tensors increases (e.g., nell-1, amazon, and reddit tensors), the likelihood of finding dense structures in the multi-dimensional space significantly decreases, leading to extremely small numbers of nonzero elements per block. Therefore, efficient execution of TD algorithms on modern parallel processors is challenging because of their low arithmetic intensity [17], random memory access, workload imbalance [18, 19], and data dependencies [17, 19]. Moreover, TD algorithms require computations along every mode, and realizing acceptable performance across all modes is difficult without using multiple mode-specific tensor copies.
@@
Prior work on the CP-ALS algorithm improved the parallel performance of the matricized tensor times Khatri-Rao product (MTTKRP) kernel [17, 18, 15, 20, 21], which is the main performance bottleneck of the overall algorithm [20]. The few performance studies conducted on the CP-APR algorithm focused primarily on performance portability [22] and streaming analysis [23], rather than parallel performance optimization. To the best of our knowledge, this paper presents the first in-depth analysis of the key performance bottlenecks of CP-APR, and significantly improves the parallel performance of both CP-ALS and CP-APR over prior state-of-the-art approaches by using a linearized mode-agnostic sparse tensor representation.
@@
Additionally, the previous approaches relied on extending legacy sparse linear algebra formats and algorithms to tensor (multilinear) algebra problems [11, 24, 20, 25, 21, 26, 15, 27, 28, 29]. These techniques can be classified based on their compression of the nonzero elements into raw or compressed formats [30, 31]. Raw formats use simple list-based representations to keep the nonzero elements along with their multi-dimensional coordinates [32]. Hence, they are mode-agnostic and typically require one tensor copy to execute tensor operations along different modes. As a result, the list-based coordinate (COO) format remains the de facto data structure for storing sparse tensors [32] in many tensor libraries (e.g., Tensor Toolbox [11], Tensorflow [33], and Tensorlab [34]). However, due to their unprocessed nature, list-based formats suffer from significant parallel and synchronization overheads [21].
@@
Compressed tensor formats [20, 35, 15, 27] use tree- or block-based structures to organize the nonzero elements, which may decrease the memory footprint of sparse tensors. However, since these approaches rely on finding clusters of nonzero elements in non-overlapping regions of the multi-dimensional space to achieve compression, their efficacy depends on the spatial data distribution, which can be highly irregular and extremely sparse as demonstrated in Figure 1. Therefore, instead of reducing memory storage, compressed formats can introduce substantial memory overhead and degrade the parallel performance of TD algorithms [30].
@@
The most popular compressed format for TD algorithms is compressed sparse fiber (CSF) [20], which extends the classical compressed sparse row (CSR) format to higher-order tensors using tree-like structures. However, CSF-based formats are mode-specific, where the arrangement of nonzero elements depends not only on the mode considered as the root of the index tree but also on the other modes at each subsequent tree level; therefore, they are typically efficient for only that specific mode order. This leads to a trade-off between performance and memory, as storing multiple tensor copies, each arranged for a specific mode, yields the best performance at the cost of extra memory [35]. For large-scale tensors, keeping multiple copies may be infeasible, especially for hardware accelerators with limited memory capacity (e.g., GPUs). Although current GPUs (e.g., H100) can have up to 80 GB of device memory [36], encoding large-scale tensors (e.g., amazon, patents and reddit tensors) using CSF requires hundreds of gigabytes of memory. In contrast, keeping only one tensor copy, arranged for an arbitrarily chosen mode order, yields the smallest memory footprint at the cost of sub-optimal performance [25]. Alternatively, memoization schemes [37, 38] utilize CSF-based formats to reduce computations by keeping and reusing intermediate results across tensor modes, which require substantial storage that largely exceeds the memory consumption of tensor representations [38]. While block-based formats [15, 27] can be mode-agnostic, their storage requirements and parallel performance still depend on the spatial data distributions as well as the parameters of the blocking/tilling schemes [27, 30], which are difficult to determine dynamically.
@@
To overcome these limitations, we present the Adaptive Linearized Tensor Order (ALTO) format. ALTO is a mode-agnostic representation that maps a set of -dimensional coordinates onto a single linearized index such that neighboring nonzero elements in the multi-dimensional space are close to each other in memory. This leads to a more cache friendly and memory-scalable tensor storage; that is, ALTO utilizes the inherent data locality of sparse tensors and its storage scales with mode lengths, rather than the number of modes. Additionally, ALTO enables a unified implementation of tensor algorithms that requires a single tensor copy to compute along all modes.
@@
In contrast to prior compressed [20, 15] and linearized [39] TD approaches, we propose a set of parallel algorithms that leverage the ALTO format to address the performance bottlenecks that have traditionally limited the scalability of sparse tensor computations. Our ALTO-based algorithms generate perfectly balanced tensor partitions in terms of the number of nonzero elements; however, these partitions may divide the multi-dimensional space of a tensor into overlapping regions (but each nonzero element belongs to a single partition). Thus, we devise data-aware adaptation heuristics that greatly improve parallel performance by locating the overlapping space between tensor partitions and selecting the best tensor traversal and conflict resolution method according to the inherent data locality of sparse tensors. Moreover, these heuristics choose between recomputing or reusing intermediate results, depending on the properties of the target tensor computations, to maximize the performance while reducing the overall memory footprint. As a result, our ALTO-based TD algorithms deliver substantial performance gains over prior state-of-the-art approaches, while allowing the processing of large-scale tensors. In summary, we make the following contributions:
@@
We present ALTO, a novel sparse tensor format for high-performance tensor algorithms. Unlike prior compressed formats, ALTO uses a single (mode-agnostic) tensor representation that improves data locality, eliminates workload imbalance, and reduces memory usage, regardless of the data distribution in the multi-dimensional space (3).
@@
We propose efficient ALTO-based parallel algorithms for the CP-ALS and CP-APR methods as well as input-aware adaptation heuristics to balance data reuse and memory footprint while greatly reducing synchronization cost (4).
@@
We conduct an in-depth performance analysis of the main TD kernels and compare against prior state-of-the-art across two generations of the latest Intel Xeon Scalable processors. The results show that on an Intel Sapphire Rapids server, ALTO-based TD algorithms achieve and geometric mean speedup over the best mode-agnostic and mode-specific formats, respectively. Compared to the best memoization method that trades lower computation for higher memory usage, ALTO attains geometric mean speedup. Furthermore, ALTO requires a fraction ( to ) of the storage used by the state-of-the-art mode-specific and memoization approaches (5).
@@ 2.1 Tensor Notations
Tensors are -dimensional arrays, where each element has an -tuple index . Each index coordinate locates a tensor element along the dimension or mode, with and . Low-dimensional tensors include vectors, where , and matrices, where . In general, a dense -dimensional (or a mode-) tensor has indexed elements. A tensor is said to be sparse if the majority of its elements are zero. The following notations are used in this paper:
@@ 2.2.1 CP-ALS
Algorithm 1 illustrates the CP-ALS algorithm for iteratively computing the factor matrices of the CPD model. In each CP-ALS iteration, every factor matrix is updated via the alternating least squares (ALS) method whereby every other factor matrix but the one being updated is fixed to yield the best approximation of . Line 13 shows the matricized tensor times Khatri-Rao product (MTTKRP) operation [20], which involves tensor matricization and Khatri-Rao product. For a mode- tensor , the mode-1 MTTKRP operation can be expressed as . MTTKRP is typically the most expensive tensor kernel of CP-ALS, and it is performed along all modes in every CP-ALS iteration. Since MTTKRP operations are similar across all modes, for brevity, we only discuss mode- MTTKRP in this paper.
@@ 2.2.2 CP-APR
While CP-ALS can decompose sparse count data, CP-APR better describes the random variations in the data by representing it using a Poisson distribution [14], which considers a discrete number of events and assumes zero probability for observing fewer than zero events. As a result, CP-APR is more expensive to compute compared to CP-ALS [14]. There are three common methods for computing CP-APR: (i) Multiplicative update (MU), (ii) Projected damped Newton for row-based sub-problems (PDN-R), and (iii) Projected quasi-Newton for row-based sub-problems (PQN-R).
@@
Algorithm 2 shows CP-APR with the MU method. Using two nested loops, CP-APR computes the decomposition of a tensor by successively updating each factor matrix while holding the other factor matrices fixed. Lines 8 and 10 show the (Khatri-Rao product) and (model update) kernels, respectively, which make up majority of the CP-APR execution time. Note that is calculated once per outer loop for each mode, and is calculated once per inner loop. Since the inner loop is executed at most times per mode per outer loop, the cost of calculating can more easily be amortized as the number of inner iterations goes up.
@@ 2.3.1 Coordinate (COO)
Decomposing a sparse tensor stored in the COO format typically involves iterating over each nonzero element and updating the corresponding factor matrix row. For example, during mode- computation, a nonzero element with coordinates (,,) updates row of the mode- factor matrix after reading rows and from mode- and mode- factor matrices, respectively. Since multiple threads can simultaneously update the same row of a factor matrix, these updates must be done atomically, which can be expensive on parallel processors with a large number of threads.